Granger Tees is a manufacturer of T-shirts. The standard amount of 100% cotton jersey fabric used to make each T-shirt is 3.00 yards. The standard price per yard of the fabric is $2.31. During last month, Granger purchased 6,300 yards of the jersey fabric for $13,860. The production department used 4,900 yards to produce 1,800 T-shirts during the month.

Compute the material price and quantity variances. (Round answers to 0 decimal places, e.g. 125. Enter all variances as a positive number.)

Material Price Variance $

FavorableNeither Unfavorable nor FavorableUnfavorable

Material Quantity Variance $

FavorableNeither Unfavorable nor FavorableUnfavorable
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (2.31 - 13,860/6300) * 6300

= 693 Favourable

.

Materials Quantity Variance = (SQ-AQ) * SP

= (1800*3 - 4900)*2.31

= 1155 Favourable

Comment if you face any issues